OpenStudy (anonymous):

this one stumps me all the time please help me What value of q makes the statement true? –(q + 14) = –3 + (–14) A. –17 B. –3 C. 3 D. 17 will give medal

OpenStudy (anonymous):

TImes both sides by -1... q+14= 3+14... q+14=17. subtract 14 from both sides. q=3

OpenStudy (anonymous):

so it would be C

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

ok i got another hold on